# Understanding the Limits of Deep Tabular Methods with Temporal Shift

### v1

@@ 4.4 Our Temporal Split
The performance and stability comparison between the random split and our newly proposed temporal split is presented in Table 1. The results demonstrate that our splitting strategy achieves a performance improvement comparable to the random split (2.18% vs. 2.17% on average) relative to the baseline temporal split in Rubachev et al. (2025). Additionally, while our temporal split shows a modest increase in the standard deviation of performance scores (AUC for classification and RMSE for regression) by 16.69%, the random split results in a much larger increase of 153.81%. This indicates that, while both methods achieve similar performance gains, our temporal split significantly outperforms the random split in terms of stability. A comparison of stability using the robustness score is provided in Appendix C.
@@ 5 What is Lost in Temporal Training?
Looking back to the MMD heatmap in Figure 4 right, we observe that the original data distribution offers a rich source of temporal information, including the periodicity and the trend. Figure 5 left presents a detailed MMD heatmap visualization of the Homesite Insurance dataset, revealing both yearly and weekly periodicity, as well as the underlying trend. We further investigate the impact of temporal shifts on deep tabular methods from the perspective of feature representations.
@@
This phenomenon may account for the suboptimal performance of datasets with clear periodic patterns, such as Cooking Time, Delivery ETA, and Maps Routing datasets, which are socially related and exhibit distinct weekly cycles, while the Weather dataset, being influenced by natural cycles, shows a clear yearly pattern. We would expect models using temporal splits to capture long-term periodicity and trends, enabling them to learn extrapolative knowledge. However, this critical knowledge does not seem to be effectively learned. In contrast, random splits appear more proficient at capturing local patterns, particularly those associated with short-term periodicity. This could explain why temporal splitting does not consistently outperform random splitting in these cases. Furthermore, it also explains the reasons why existing methods encounter challenges in dealing with temporal shifts.
@@ 6 Temporal Embeddings
Building on the above analysis, it is essential to incorporate temporal information into the model in a manner that effectively captures the underlying temporal dependencies. To address this issue, we propose a lightweight, plug-and-play temporal embedding method specifically designed for timestamps, aiming to investigate whether providing explicit temporal information through embedding can lead to performance improvements. Empirically, timestamps are often treated as noise and discarded. However, in the context of temporal shifts and temporal splitting, timestamps likely contain crucial temporal information that enables the model to align with the periodicity and trends inherent in the temporal sequence. Existing works have also emphasized that, in certain contexts, timestamps serve as a valuable feature (Wang et al., 2024; Zeng et al., 2024).
@@
In our embedding method, We fit multi-scale periodicity using Fourier series expansion (Tancik et al., 2020; Li et al., 2021). The Fourier series can be expressed as:
@@
where is the timestamp, and the two components of the embedding each capture the periodicity and trend of the timestamp, respectively. Those are further defined as
where , are given periodicity priors for Fourier-based embedding. In our experiments, we set for yearly, monthly, weekly, and daily periodicity. These periodicity are chosen based on the common temporal patterns observed in the datasets, which effectively capture natural patterns in datasets with inherent temporal correlations, such as Weather, by modeling yearly and daily periodicity, and societal-related temporal patterns, such as monthly and weekly periodicity, in datasets like Homesite Insurance and Cooking Time. Each Fourier embedding is defined as
for , which represents the process of Fourier series expansion. By decomposing multiple periods into the components of a Fourier series and passing them through a learnable linear layer, the process of solving and aggregating parameters and in the Fourier series can be simulated, resulting in a set of periodic functions that incorporate multiple cycles. The ReLU activation further facilitates the selection process, ensuring the quality of the learned periodic information.
@@
PLR: Applying a learnable one-dimension PLR embedding (Gorishniy et al., 2022) to the timestamp, enabling the model to capture temporal patterns adaptively.
@@
The models’ performance improvement after adopting our temporal embedding are presented in Figure 7. Most methods demonstrate improvements, highlighting the importance of leveraging temporal information in addressing temporal shift tasks. Notably, MLP, DCNv2, and FT-T show significant improvements, while as previously mentioned, MLP-PLR, TabR, TabM, and ModernNCA exhibit limited gains, likely due to their incompatibility with numerical embeddings. This suggests that temporal features may require dedicated embedding strategies rather than relying on existing numerical embedding approaches. After applying temporal embedding, both ModernNCA and TabR demonstrate strong performance, indicating that with an appropriate training protocol and temporal embedding, even retrieval-based methods, which are typically most affected by distributional shift, can regain their practical utility.
The MMD heatmap of the model representation after adopting temporal embedding is shown in Figure 5 right. The patterns are closer to the original data, reflecting that it captures rich and correct temporal information, thus effectively alleviating the loss of temporal information during training. The reappearance of diagonal stripes indicates that the model has learned independent representations for each temporal phase within the period, thereby confirming the adaptive role of temporal embedding.
@@ Appendix A Dataset Explanation
The detailed information about the dataset is provided in Table 3. We used the dataset from TabReD (Rubachev et al., 2025) in its entirety and applied the same preprocessing method.
@@
In Figure 3, Table 1, and Table 2, we compare the average performance improvement for each method under different strategies (e.g. splitting or temporal embedding), specifically the percentage increase in AUC on classification datasets and the percentage decrease in RMSE on regression datasets. In Figure 2 and Figure 7, we present a comparison of performance across different training protocols and methods, where all results are reported as performance improvements relative to the MLP performance under the original split in (Rubachev et al., 2025). Since performance improvements between methods can often be influenced by outliers, we apply a robust average, excluding the maximum and minimum performance improvements across the eight datasets before calculating the mean.
@@
The only distinction in the implementation is that TabReD employs different embedding methods for numerical and categorical features for each method-dataset pair. Specifically, it uses identity or noisy-quantile encoding for numerical features and one-hot or ordinal encoding for categorical features. To ensure a fair comparison, we reproduced the experiment from TabReD by removing numerical embeddings and fixing categorical embeddings to one-hot encoding where necessary. This adjustment is essential for accurately evaluating the performance of the methods and for advancing future research on temporal embeddings.
@@
It is important to note that the HomeCredit Default dataset suffers from severe class imbalance, which makes it challenging for methods with limited feature extraction capabilities, such as MLP, SNN (Klambauer et al., 2017), and DCNv2 (Wang et al., 2021b), to perform well without additional numerical feature embeddings. While the AUC of the naive MLP (as well as SNN and DCNv2 methods) on this dataset drops to around 0.55, rendering the comparison with other methods (which typically achieve an AUC greater than 0.80) meaningless, it is noteworthy that the insights on training protocols and temporal embeddings explored in this work still lead to significant performance improvements on this dataset, as shown in Appendix C.
@@
At the same time, our experiment on the equivalence (Figure 3) of the validation set shows that the datasets whose experimental results seriously fail to meet the equivalence assumption, such as Sberbank Housing and Ecom Offers, are confirmed to have the most serious nonuniformity in sampling. The temporal sampling distribution for each dataset is shown in Figure 8. Instead of focusing on the time span between the training and validation sets, we concentrate on ensuring that the number of instances in both sets is equal across different splits, thereby enabling a fair comparison of the splitting strategies. When the temporal sampling distribution becomes highly non-uniform, the observation in Figure 4 no longer holds, as it relies on a constant time span for the validation set. Consequently, our insights regarding the equivalence of the validation sets are significantly affected in such situations.
This indicates that the conclusion regarding the equivalence of the validation set on different temporal direction is actually stronger, especially for datasets with uniform sampling over time.
@@ Appendix B Experimental Setup
We adopt training, evaluation and tuning setup from Ye et al. (2024) and Liu et al. (2024). We tune hyper-parameters using Optuna (Akiba et al., 2019), performing 100 trials for most methods to identify the best configuration. The hyper-parameter search space follows exactly the settings in Rubachev et al. (2025). Using these optimal hyper-parameters, each method is trained with 15 random seeds, and the average performance across seeds is reported. We label the attribute type (numerical or categorical) for gradient boosting methods such as CatBoost. For all deep learning methods, we use a batch size of 1024 and AdamW (Loshchilov & Hutter, 2019) as the optimizer.
We followed Rubachev et al. (2025) and performed only 25 hyper-parameter tuning runs for FT-T and TabR, as these methods exhibit lower efficiency on datasets with large feature dimensions and sample sizes. For our temporal embedding, we conducted separate hyper-parameter searches for the periodic order and linear trend. However, since 25 tuning trials
For classification tasks, we evaluate models using AUC (higher is better) as the primary metric and use RMSE (lower is better) for regression tasks to select the best-performing model during training on the validation set.
To ensure the validity of random splitting, each group of random split experiments was tested on three distinct random splits, with 15 random seeds run on each split. The mean performance across these runs is reported as the final result. The variance of the random split is calculated based on all 45 results (3 splits × 15 seeds), as the random split is subject to variance from both the split selection and the running seeds during the training phase. This approach better reflects the overall stability of the standard procedure.
@@ C.1 Splitting Strategies
We present the detailed results of our experiments with different splitting strategies in Table 4. Comparisons between the baseline temporal split from Rubachev et al. (2025), the random split discussed in Section 1, and our proposed temporal split in Section 4.4 are shown in Table 5 and Table 6. Additionally, we introduce a new experiment that highlights the stability of our splitting strategy using the robustness score, as illustrated in Figure 9.
@@
The thorough comparison of the original temporal split in (Rubachev et al., 2025), the random split in Section 1, and our newly proposed temporal split in Section 4.4 is demonstrated in Figure 9. Our proposed splitting strategy results in a comparable performance to random split while offering better stability. Since random split suffers a severe instability shown in Table 1, including the splitting selection and the running seed, which cannot be estimated at the training stage, we use robustness score for comparison, which is defined as
@@ C.2 Temporal Embeddings
We present the detailed performance results of different embedding methods discussed in Section 6 in Table 7, as well as the model performance before and after adopting our proposed temporal embedding in Table 8.

### v2

@@ 4.4 Our Temporal Split
The performance and stability comparison between the random split and our newly proposed temporal split is presented in Table 1. The results demonstrate that our splitting strategy achieves a performance improvement comparable to the random split (2.18% vs. 2.17% on average) relative to the baseline temporal split in rubachev2024tabred. Additionally, while our temporal split shows a modest increase in the standard deviation of performance scores (AUC for classification and RMSE for regression) by 16.69%, the random split results in a much larger increase of 153.81%. This indicates that, while both methods achieve similar performance gains, our temporal split significantly outperforms the random split in terms of stability. A comparison of stability using the robustness score is provided in Section C.1.
@@ 5 What is Lost in Temporal Training?
Looking back to the MMD heatmap in Figure 4 right, we observe that the original data distribution offers a rich source of temporal information, including the periodicity and the trend. Figure 5 left presents a detailed MMD heatmap visualization of the HI dataset, revealing both yearly and weekly periodicity, as well as the underlying trend. We further investigate the impact of temporal shifts on deep tabular methods from the perspective of feature representations.
@@
This phenomenon may account for the suboptimal performance of datasets with clear periodic patterns, such as CT, DE, and MR datasets, which are socially related and exhibit distinct weekly cycles, while the WE dataset, being influenced by natural cycles, shows a clear yearly pattern. We would expect models using temporal splits to capture long-term periodicity and trends, enabling them to learn extrapolative knowledge. However, this critical knowledge does not seem to be effectively learned. In contrast, random splits appear more proficient at capturing local patterns, particularly those associated with short-term periodicity. This could explain why temporal splitting does not consistently outperform random splitting in these cases. Furthermore, it also explains the reasons why existing methods encounter challenges in dealing with temporal shifts.
@@ 6 Temporal Embeddings
Building on the above analysis, it is essential to incorporate temporal information into the model in a manner that effectively captures the underlying temporal dependencies. To address this issue, we propose a lightweight, plug-and-play temporal embedding method specifically designed for timestamps, aiming to investigate whether providing explicit temporal information through embedding can lead to performance improvements. Empirically, timestamps are often treated as noise and discarded. However, in the context of temporal shifts and temporal splitting, timestamps likely contain crucial temporal information that enables the model to align with the periodicity and trends inherent in the temporal sequence. Existing works have also emphasized that, in certain contexts, timestamps serve as a valuable feature (wang2024rethinking; zeng2024howmuch).
@@
In our embedding method, we fit multi-scale periodicity using Fourier series expansion (tancik2020fourier; li2021learnable). The Fourier series can be expressed as:
@@
where is the timestamp, and the two components of the embedding each capture the periodicity and trend of the timestamp. The former part is further defined as
where are given periodicity priors for Fourier-based embedding. In our experiments, we set to represent yearly, monthly, weekly, and daily periodicities. These choices reflect common temporal patterns observed in the datasets. For example, in datasets with natural temporal correlations such as WE, yearly and daily periodicities are effective, while datasets like HI and CT benefit from modeling socially driven patterns captured by monthly and weekly periodicities. Pre-defined periodicity priors yield more stable and interpretable results, as illustrated in Section C.2. Each Fourier-based embedding is defined as
for orders , representing the -term Fourier series expansion. We assign the Fourier orders individually for each periodic prior. By combining multiple periodic components and projecting them through a learnable linear layer, the embedding approximates the aggregation of Fourier coefficients and as formulated in Equation 1. A subsequent activation promotes sparsity and emphasizes salient frequencies, thereby improving the quality of the learned periodic representations.
@@
PLR: Applying PLR embedding (gorishniy2022embeddings) to the timestamp, treating the resulting representations as numerical features to capture temporal patterns adaptively.
@@
The models’ performance improvement after adopting our temporal embedding is presented in Figure 7. Most methods demonstrate improvements, highlighting the importance of leveraging temporal information in addressing temporal shift tasks. As previously mentioned, MLP-PLR and TabM exhibit limited gains, likely due to their incompatibility with PLR embedding. However, these models achieve substantially greater improvements when the temporal embedding is directly integrated into the model backbone, as shown in Section C.2. This suggests that temporal features may require dedicated embedding strategies rather than relying on existing numerical embedding approaches. After applying temporal embedding, both ModernNCA and TabR demonstrate strong performance, indicating that with an appropriate training protocol and temporal embedding, even retrieval-based methods, which are typically most affected by distributional shift, can regain their practical utility.
The MMD heatmap of the model representation after adopting temporal embedding is shown in Figure 5 right and Section C.4. The patterns are closer to the original data, reflecting that it captures correct temporal information, thus effectively alleviating the loss of temporal information during training. The reappearance of diagonal stripes indicates that the model has learned independent representations for each temporal phase within the period, thereby confirming the adaptive role of temporal embedding.
@@ 7 Conclusion
Acknowledgement
This work is partially supported by National Key R&D Program of China (2024YFE0202800), NSFC (62376118), Key Program of Jiangsu Science Foundation (BK20243012), Collaborative Innovation Center of Novel Software Technology and Industrialization. We thank Si-Yang Liu, Qi-Le Zhou, and Jun-Peng Jiang for their insightful discussions.
Impact Statement
This paper aims to advance the field of machine learning by addressing the critical challenge of temporal distribution shifts in tabular data, which frequently occur in real-world applications. The proposed training protocol and temporal embedding method offer practical improvements for deploying existing tabular models in open environments.
@@ Appendix A Dataset Explanation
The detailed information about the dataset is provided in Table 3. We adopt the full TabReD dataset (rubachev2024tabred) without modification.
@@
In Figure 3, Table 1, and Table 2, we compare the average performance improvement for each method under different strategies (e.g. splitting or temporal embedding), specifically the percentage increase in AUC on classification datasets and the percentage decrease in RMSE on regression datasets. In Figure 2 and Figure 7, we present a comparison of performance across different training protocols and methods, where all results are reported as performance improvements relative to the MLP performance under the original split in rubachev2024tabred. Since performance improvements between methods can often be influenced by outliers, we apply a robust average, excluding the maximum and minimum performance improvements across the eight datasets before calculating the mean.
@@
The only distinction in the implementation is that TabReD employs different encoding methods during preprocessing for numerical and categorical features for each method-dataset pair. Specifically, it uses identity or noisy-quantile encoding for numerical features and one-hot or ordinal encoding for categorical features. To ensure a fair comparison, we reproduced the experiment from TabReD by removing numerical encoding and fixing categorical encoding to one-hot encoding where necessary. This adjustment is essential for accurately evaluating the performance of the methods and for advancing future research on temporal embeddings.
@@
It is also important to note that the HD dataset suffers from severe class imbalance, which makes it challenging for methods with limited feature extraction capabilities, such as MLP, SNN (klambauer2017self), and DCNv2 (wang2021dcn), to perform well without additional numerical feature encoding. On this dataset, the AUC of the naive MLP (as well as SNN and DCNv2 methods) drops to approximately 0.55, in contrast to other methods that typically achieve an AUC above 0.80. This large discrepancy renders direct comparison uninformative. To mitigate this issue, we adopt a robust average when reporting the percentage improvement over MLP. Nevertheless, the insights on training protocols and temporal embeddings explored in this work still lead to significant performance improvements on this challenging dataset, as detailed in Appendix C.
@@
At the same time, our experiment on the equivalence (Figure 3) of the validation set shows that the datasets whose experimental results seriously fail to meet the equivalence assumption, such as SH and EO, are confirmed to have the most serious nonuniformity in sampling. The temporal sampling distribution for each dataset is shown in Figure 8.
Instead of focusing on the time span between the training and validation sets, we concentrate on ensuring that the number of instances in both sets is equal across different splits, thereby enabling a fair comparison of the splitting strategies. When the temporal sampling distribution becomes highly non-uniform, the observation in Figure 4 no longer holds, as it relies on a constant time span for the validation set. Consequently, our insights regarding the equivalence of the validation sets are significantly affected in such situations. This indicates that the conclusion regarding the equivalence of the validation set on different temporal directions is actually stronger, especially for datasets with uniform sampling over time.
@@ Appendix B Experimental Setup
Our source code is now available at https://github.com/LAMDA-Tabular/Tabular-Temporal-Shift.
We adopt preprocessing, training, evaluation and tuning setup from ye2024closer and liu2024talent. We tune hyper-parameters using Optuna (akiba2019optuna), performing 100 trials for most methods (except FT-T and TabR for only 25, as explained below) to identify the best configuration. The hyper-parameter search space follows exactly the settings in rubachev2024tabred, and can be referred from our source code. The hyper-parameter search space of our proposed temporal embedding is summarized separately in Table 4. Using these optimal hyper-parameters, each method is trained with 15 random seeds, and the average performance across seeds is reported. We label the attribute type (numerical or categorical) for gradient boosting methods such as CatBoost. For all deep learning methods, we use a batch size of 1024 and AdamW (loshchilov2019decoupled) as the optimizer, with an early stopping patience of 16. For classification tasks, we evaluate models using AUC (higher is better) as the primary metric and use RMSE (lower is better) for regression tasks to select the best-performing model during training on the validation set.
| Parameter | Distribution |
| year_order | |
| month_order | |
| day_order | |
| hour_order | |
| trend | |
| d_embedding |
We followed rubachev2024tabred and performed only 25 hyper-parameter tuning trials for FT-T and TabR, as these methods exhibit lower efficiency on datasets with large feature dimensions and sample sizes. For our temporal embedding, we conducted separate hyper-parameter searches for the periodic order and linear trend, to selectively extract temporal information while reducing computational overhead. However, since 25 tuning trials were insufficient to identify the optimal hyper-parameters for the temporal embedding, we performed a global tuning of the temporal embedding order for FT-T and TabR, with all periodic components sharing the same order.
To ensure the validity of random splitting, each group of random split experiments was tested on three distinct random splits, with 15 random seeds run on each split. The mean performance across these runs is reported as the final result. The standard deviation of the random split is calculated based on all 45 results (3 splits × 15 seeds), as the random split is subject to variability from both the split selection and the running seeds during the training phase. This approach better reflects the overall stability of the standard procedure.
@@ C.1 Splitting Strategies
| Methods | Splits | HI | EO | HD | SH | CT | DE | MR | WE | Avg. Imp. |
| Mambular | Original | 0.6782 | 0.5713 | 0.5872 | 0.6822 | 0.5107 | 0.5981 | 0.2330 | 2.2232 | – |
| Ours | 0.6870 | 0.6010 | 0.5396 | 0.5492 | 0.5080 | 0.5999 | 0.2297 | 2.1972 | 2.59% | |
| TabPFN v2 | Original | 0.8215 | 0.5660 | 0.5000 | 0.2276 | 0.4858 | 0.5529 | 0.1673 | 1.7076 | – |
| Ours | 0.8205 | 0.5717 | 0.5000 | 0.2204 | 0.4804 | 0.5619 | 0.1740 | 1.6031 | 0.71% |
To further validate the general applicability of our training protocol, we apply our temporal splitting strategy to two new model families: the autoregressive model Mambular (thielmann2024mambular) and the tabular general model TabPFN-v2 (hollmann2025accurate). As shown in Table 5, our temporal split consistently improves performance compared to the original splits used by these models.
We further report detailed experimental results for the different data splitting strategies illustrated in Figure 3, with corresponding metrics presented in Table 6. Specifically, Table 7 and Table 8 compare the performance and standard deviation across three strategies: the baseline temporal split from rubachev2024tabred, the random split discussed in Section 1, and our proposed temporal split described in Section 4.4. These results demonstrate the superior performance and stability of our refined temporal split. Furthermore, Figure 9 presents an additional experiment using the robustness score to further highlight the consistency of our splitting strategy.
@@
The thorough comparison of the original temporal split in (rubachev2024tabred), the random split in Section 1, and our newly proposed temporal split in Section 4.4 is demonstrated in Figure 9. Our proposed splitting strategy results in a comparable performance to random split while offering better stability. Since random split suffers a severe instability shown in Table 1, including the splitting selection and the running seed, which cannot be estimated at the training stage, we use robustness score for comparison, which is defined as
@@ C.2 Temporal Embeddings
We conducted a further analysis on the incompatibility between temporal embedding and PLR embedding discussed in Section 6. Our temporal embedding converts timestamps into numerical inputs, which may be incompatible with PLR numerical embedding (gorishniy2022embeddings). Specifically, once timestamps are embedded, their representation reflects temporal similarity. Applying another periodic transformation via PLR could increase optimization difficulty. Directly feed the temporal embedding into the model backbone consistently improves, as demonstrated in Table 9.
| Methods | Emb. | HI | EO | HD | SH | CT | DE | MR | WE | Avg. Imp. |
| Ours⋆ | 0.9609 | 0.6215 | 0.8155 | 0.2337 | 0.4788 | 0.5492 | 0.1616 | 1.5152 | 0.26% | |
| Ours | 0.9629 | 0.6271 | 0.8363 | 0.2321 | 0.4791 | 0.5488 | 0.1609 | 1.4812 | 0.07% | |
| Ours⋆ | 0.9633 | 0.6260 | 0.8368 | 0.2297 | 0.4775 | 0.5490 | 0.1607 | 1.4873 | 0.20% | |
| Ours⋆ | 0.9623 | 0.6357 | 0.8344 | 0.2262 | 0.4790 | 0.5515 | 0.1616 | 1.4855 | 0.41% |
Feed temporal embedding directly to the backbone.
We also experimented with tuning hyperparameters to find the cycles, results in Table 10. When using fixed prior periods, ModernNCA achieved a 0.30% performance improvement, while setting adjustable cycles resulted in a -2.48% performance decline. In temporal distribution shift scenarios, due to the absence of an entirely accurate validation set, we believe that prior knowledge of fixed cycles is more stable and interpretable than adjustable cycles. It’s also important to note that in many tasks, complete cycles are not available. For example, in the WE dataset, there is a yearly cycle, but the training set does not span a full year, which highlights the importance of prior knowledge.
| Methods | Cycles | HI | EO | HD | SH | CT | DE | MR | WE | Avg. Imp. |
| Fixed | 0.9620 | 0.6356 | 0.8316 | 0.2255 | 0.4791 | 0.5535 | 0.1617 | 1.4903 | 0.30% | |
| Adjustable | 0.9575 | 0.5719 | 0.8355 | 0.2469 | 0.4820 | 0.5493 | 0.1630 | 1.4926 | 2.20% |
We further present the detailed performance results of different embedding methods discussed in Section 6 in Table 11, as well as the model performance before and after adopting our proposed temporal embedding in Table 12.
@@ C.3 Performance Rankings
It is worth noting that although the relative improvement of TabM over MLP decreases after adding the temporal embedding (0.12% in Figure 7), TabM itself still achieves a 0.07% performance gain (shown in Table 12) and rises by 0.875 ranks in the average model ranking (shown in Table 13). Notably, none of the other methods experience a performance drop. This provides a comprehensive multi-perspective evaluation.
@@
C.4 Model Representations
We present the model representations of TabM before and after incorporating temporal embedding in Figure 10, as a representative SOTA model complementary to the MLP shown in Figure 5. Since our TabM internally ensembles 32 model representations, we compute their average to reduce computational cost, which aligns with how the model operates during inference. The left plot exhibits a clear grid structure and shows richer representations than MLP, reflecting the stronger representational capacity of TabM. However, the absence of diagonal stripes suggests that the model still primarily captures only two coarse temporal modes (i.e., weekdays and weekends). In contrast, the right plot shows the model representation after incorporating temporal embedding, where both the grid structure and diagonal patterns are clearly visible, indicating that the model has learned more fine-grained temporal knowledge.